# Congruence subgroup property for nilpotent groups and subsurface subgroups of Mapping Class Groups

### v1

@@ Abstract.
We prove the Congruence Subgroup Property for two families of subgroups of Mapping Class Groups of finite-type surfaces, the first one related to nilpotent quotients of the fundamental group and Johnson filtration, and the second one consisting of certain geometric subgroups obtained via subsurface inclusions.
@@ 1. Introduction
Groups possessing the Congruence Subgroup Property include $\textnormal{SL}_{n}\mathbb{Z}$ for $n\geqslant 3$, by on the work of Mennicke [Men65] and Bass–Lazard–Serre [BMS67] (see also [PR10]). Another groups with CSP are $\textnormal{Aut}(F_{2})$ and $\textnormal{Out}(F_{2})$, by the work of Asada [Asa01][BER11]. An example where CSP fails is $\textnormal{SL}_{2}\mathbb{Z}$ (this is one of the consequences of its virtual freeness). This is somewhat surprising, given that the natural map $\textnormal{Out}(F_{2})\to\textnormal{GL}_{2}\mathbb{Z}$ is an isomorphism; the explanation is that the actions of this group on $F_{2}$ and $\mathbb{Z}^{2}$ induce different families of congruence subgroups.
@@
This article contains two main results. The first one, Theorem 4, holds for all nilpotent groups, and says that CSP can be determined by the action on abelianisation. We use it to show that finite-index subgroups of Mapping Class Groups containing some term of the Johnson filtration are congruence subgroups. The second, Theorem 7, asserts that CSP of Mapping Class Groups behaves well under gluing smaller surfaces along their boundaries. As corollaries we show that many natural geometric subgroups of Mapping Class Groups have the Congruence Subgroup Property.
@@ 1.1. Outer automorphisms of nilpotent groups
The first result of this article is that outer automorphism groups of nilpotent groups posess the congruence subgroup property, as long as their abelianisations do.
@@
Theorem 4 is proved in Section 4.2.
@@ 1.2. Subsurface subgroups of MCGs
Let $\Sigma$ be a possibly punctured surface with no boundary, and let $\Sigma=\bigcup_{i}\Sigma_{i}$ be its decomposition into disjoint subsurfaces glued along boundaries. Suppose that $\Lambda_{i}\leq\textnormal{Mcg}(\Sigma_{i})$ are subgroups suported on the subsurfaces, and let $\Lambda=\prod_{i}\Lambda_{i}$ be the subgroup of $\textnormal{Mcg}(\Sigma)$ obtained from them by inclusion. If all $\Lambda_{i}\to\textnormal{Out}(\pi_{1}\Sigma_{i})$ have the CSP, then $\Lambda\to\textnormal{Out}(\pi_{1}\Sigma)$ has the CSP.
@@ Acknowledgements
The author would like to thank Dawid Kielak for suggesting this project and for his support throughout. This work was supported by the Simons Investigator Award 409745 from the Simons Foundation.
@@ 4. CSP for automorphisms of nilpotent groups
The goal of this section is to prove Theorem 4. We start by proving in Subsection 4.1 two Lemmata 18 and 19, which are technical tools for producing special quotients of nilpotent groups. We complete the proof of Theorem 4 at the end of Subsection 4.2, after proving its special case Proposition 20.
@@ 4.2. Proof of Theorem 4
As a step towards Theorem 4 we prove Proposition 20, which gives us CSP of arbitary subgroups of the last term of the Andreadakis–Johnson filtration. It follows from Lemmata 18 and 19.
@@
We know from Lemma 16 that $\delta\colon\mathcal{I}^{(K)}\to D_{K}$ is injective. Moreover, $D_{K}$ is a finitely generated $\mathbb{Z}$-module. Therefore, for every finite-index $\Gamma\leq\Lambda$ there exists $M\in\mathbb{N}_{+}$ such that $\delta(\Lambda)\cap M\cdot D_{K}\subseteq\delta(\Gamma)$. To finish the proof it suffices to show that for every $M\in\mathbb{N}_{+}$ there exists a finite characteristic quotient $r\colon G\twoheadrightarrow R$ such that
@@ 5.3. Proof of Theorem 7
We are ready to complete the proof of Theorem 7. This involves combining Proposition 23 with Lemmata 21 and 13. We recall the statement for the reader’s convenience.
Theorem (restatement of 7).
Let $\Sigma$ be a possibly punctured surface with no boundary, and let $\Sigma=\bigcup_{i}\Sigma_{i}$ be its decomposition into disjoint subsurfaces glued along boundaries. Suppose that $\Lambda_{i}\leq\textnormal{Mcg}(\Sigma_{i})$ are subgroups suported on the subsurfaces, and let $\Lambda=\prod_{i}\Lambda_{i}$ be the subgroup of $\textnormal{Mcg}(\Sigma)$ obtained from them by inclusion. If all $\Lambda_{i}\to\textnormal{Out}(\pi_{1}\Sigma_{i})$ have the CSP, then $\Lambda\to\textnormal{Out}(\pi_{1}\Sigma)$ has the CSP.
@@ 6. Example application: handle pushing subgroups
Additionally, the point-pushing subgroup is precisely the kernel of the homomorphism $\textnormal{Mcg}(\Sigma_{g,p,b})\to\textnormal{Mcg}(\Sigma_{g,0,0})$ obtained by filling the punctures and capping the boundary components with discs. This statement is known as the Birman exact sequence [Bir69, Theorem 1], [FM12, Theorem 4.6].

### v2

@@ Abstract.
We prove the Congruence Subgroup Property for two families of subgroups of Mapping Class Groups of finite-type surfaces. The first one is related to nilpotent quotients of the fundamental group and Johnson filtration, and along the way we give an elementary proof of a theorem of Ben-Ezra–Lubotzky. The second family consists of certain geometric subgroups obtained via subsurface inclusions.
@@ 1. Introduction
Groups possessing the Congruence Subgroup Property include $\textnormal{SL}_{n}\mathbb{Z}$ for $n\geqslant 3$, by the work of Mennicke [Men65] and Bass–Lazard–Milnor–Serre [BLS64][BMS67] (see also [PR10]). Another two groups with CSP are $\textnormal{Aut}(F_{2})$ and $\textnormal{Out}(F_{2})$, by the work of Asada [Asa01] (see also [BER11][BEL18]). An example where CSP fails is $\textnormal{SL}_{2}\mathbb{Z}$ (this is one of the consequences of its virtual freeness). This is somewhat surprising, given that the natural map $\textnormal{Out}(F_{2})\to\textnormal{GL}_{2}\mathbb{Z}$ is an isomorphism; the explanation is that the actions of this group on $F_{2}$ and $\mathbb{Z}^{2}$ induce different families of congruence subgroups.
@@
This article contains two main theorems. The first one, Theorem 4, holds for all nilpotent groups, and says that CSP can be determined by the action on abelianisation. This result is originally due to Ben-Ezra–Lubotzky [BEL22]; we present another proof that is elementary and requires no background theory. We use Theorem 4 to show that finite-index subgroups of Mapping Class Groups containing some term of the Johnson filtration are congruence subgroups. The second result, Theorem 8, asserts that CSP of Mapping Class Groups behaves well under gluing smaller surfaces along their boundaries. As corollaries we show that many natural geometric subgroups of Mapping Class Groups have the Congruence Subgroup Property.
@@ 1.1. Outer automorphisms of nilpotent groups
The first theorem of this article is that outer automorphism groups of nilpotent groups possess the congruence subgroup property, as long as their abelianisations do.
@@
This result (up to torsion) was first obtained by Ben-Ezra and Lubotzky [BEL22, Theorem 1.1], basing on the theory of Malcev completion and unipotent groups and knowledge of CSP for solvable arithmetic groups. In Section 4.2 we give a self-contained proof of Theorem 4 involving only induction and elementary algebra.
Remark 5.
Metabelian groups behave rather differently in terms of CSP [BEL18].
@@ 1.2. Subsurface subgroups of MCGs
Let $\Sigma$ be a possibly punctured surface with no boundary, and let $\Sigma=\bigcup_{i}\Sigma_{i}$ be its decomposition into disjoint subsurfaces glued along boundaries. Suppose that $\Lambda_{i}\leq\textnormal{Mcg}(\Sigma_{i})$ are subgroups supported on the subsurfaces, and let $\Lambda=\prod_{i}\Lambda_{i}$ be the subgroup of $\textnormal{Mcg}(\Sigma)$ obtained from them by inclusion. If all $\Lambda_{i}\to\textnormal{Out}(\pi_{1}\Sigma_{i})$ have the CSP, then $\Lambda\to\textnormal{Out}(\pi_{1}\Sigma)$ has the CSP.
@@ Acknowledgements
The author is grateful to Dawid Kielak for suggesting this project and for his support throughout. He would also like to thank David El-Chai Ben-Ezra and Alex Lubotzky for valuable comments on an earlier version of this manuscript. This work was supported by the Simons Investigator Award 409745 from the Simons Foundation.
@@ 4. CSP for automorphisms of nilpotent groups
The goal of this section is to prove Theorem 4. We start by proving in Subsection 4.1 two Lemmata 19 and 20, which are technical tools for producing special quotients of nilpotent groups. We complete the proof of Theorem 4 at the end of Subsection 4.2, after proving its special case Proposition 21.
The overall proof approach is straightforward: to induct on the length of central series with the help of Lemma 14. This is perhaps closest to the strategy outlined in [BEL22, Section 4]; however, thanks to Lemmata 19 and 20, we will be able to complete the inductive step for arbitrary (including non-free) nilpotent groups.
@@ 4.2. Proof of Theorem 4
As a step towards Theorem 4 we prove Proposition 21, which gives us CSP of arbitary subgroups of the last term of the Andreadakis–Johnson filtration. It follows from Lemmata 19 and 20.
@@
Lemma 17 provides us with an injective group homomorphism $\delta\colon\mathcal{I}^{(K)}\to D_{K}$. Moreover, $D_{K}$ is a finitely generated $\mathbb{Z}$-module. Therefore, for every finite-index $\Gamma\leq\Lambda$ there exists $M\in\mathbb{N}_{+}$ such that $\delta(\Lambda)\cap M\cdot D_{K}\subseteq\delta(\Gamma)$. To finish the proof it suffices to show that for every $M\in\mathbb{N}_{+}$ there exists a finite characteristic quotient $r\colon G\twoheadrightarrow R$ such that
@@ 5.3. Proof of Theorem 8
We are ready to complete the proof of Theorem 8. This involves combining Proposition 24 with Lemmata 22 and 14. We recall the statement for the reader’s convenience.
Theorem (restatement of 8).
Let $\Sigma$ be a possibly punctured surface with no boundary, and let $\Sigma=\bigcup_{i}\Sigma_{i}$ be its decomposition into disjoint subsurfaces glued along boundaries. Suppose that $\Lambda_{i}\leq\textnormal{Mcg}(\Sigma_{i})$ are subgroups supported on the subsurfaces, and let $\Lambda=\prod_{i}\Lambda_{i}$ be the subgroup of $\textnormal{Mcg}(\Sigma)$ obtained from them by inclusion. If all $\Lambda_{i}\to\textnormal{Out}(\pi_{1}\Sigma_{i})$ have the CSP, then $\Lambda\to\textnormal{Out}(\pi_{1}\Sigma)$ has the CSP.
@@ 6. Example application: handle pushing subgroups
Additionally, the point-pushing subgroup is precisely the kernel of the homomorphism $\textnormal{Mcg}(\Sigma_{g,p,b})\to\textnormal{Mcg}(\Sigma_{g,0,0})$ obtained by filling the punctures and capping the boundary components with discs. This statement is known as the Birman exact sequence [Bir69, Theorem 1][FM12, Theorem 4.6].